The sum of three numbers is 18. The second number is six more than the first. The third is three less than the second number. What are the numbers?
Answer by rfer(16322)

x+x+6+x+6-3=18
3x+9=18
3x=18-9
3x=9
x=9/3
x=3
==============
3+3+6+3+6-3=18